Find the area and perimeter of the polygon with vertices A(-3,0), B(3,4) C(5,1),and D(-1,-3)
Answer by Alan3354(69443)

Find the area and perimeter of the polygon with vertices A(-3,0), B(3,4) C(5,1),and D(-1,-3)
Area:
-3 3 5 -1 -3
0 4 1 -3 0
--> -12 + 3 -15 + 0 = -24 down slope
--> 0 + 20 -1 + 9 = 28 up slope
Diff = 52
Area = 26 sq units
Email me if you want to see how to do that at Moral Loophole@aol.com
---------
Perimeter:
Slope of AB = slope of DC
Slope of BC = slope of AC
Slope of AB = -1/slope of AD
--> a rectangle
Length of AB = sqrt(6^2 + 4^2) = sqrt(52)
Length of AD = sqrt(2^2 + 3^2) = sqrt(13)
Perimeter = 2*(AB + AD) = 6sqrt(13)
-------------
Area = sqrt(52)*sqrt(13) = sqrt(676)
Area = 26, same answer as above